OpenStudy (anonymous):

what is 4 2/7 X 2 1/5? express in simplest form.

OpenStudy (phi):

can you write 4 2/7 as an improper fraction ?

OpenStudy (anonymous):

its 30/7

OpenStudy (phi):

yes. now do the same for 2 1/5

OpenStudy (anonymous):

11/5

OpenStudy (phi):

now you have \[ \frac{30}{7}\cdot \frac{11}{5} \] you multiply top times top and bottom times bottom BUT... it's simpler to first look for numbers that you can simplify. notice 30 can be divided by 5 and 5 (in the bottom) can be divided by 5 so do that: \[ \frac{\cancel{30}6}{7}\cdot \frac{11}{\cancel{5}1} \] now do top times top and bottom times bottom

OpenStudy (anonymous):

ok thanks :)

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

66/7

OpenStudy (phi):

now change back to a whole number plus a fraction.

OpenStudy (anonymous):

that would be 9 and something over umm7

OpenStudy (phi):

the remainder over the divisor (i.e over 7) you can check your answer by changing it back into an improper fraction .. if you did it correctly, you get back 66/7

OpenStudy (anonymous):

ya i know that the answer is 9 and something over 7 but i cant get that part cuz i dont know how

OpenStudy (phi):

7 goes into 66 9 times 7*9= ?

OpenStudy (anonymous):

63

OpenStudy (phi):

yes. Here is how you divide: